OpenStudy (anonymous):

r = <9, -7, -8>, v = <3, 4, 7>, w = <6, -9, 7> v ⋅ w

OpenStudy (danjs):

vector dot product is a scalar

OpenStudy (anonymous):

what does that mean?

OpenStudy (danjs):

\[v * w = <v _{x}, v _{y,} v _{z}> * <w _{x}, w _{y} , w _{z}>\]

OpenStudy (danjs):

\[= v _{x}w _{x} + v _{y} w _{y} + v _{z}w _{z}\]

OpenStudy (danjs):

does that look familiar?

OpenStudy (anonymous):

no @DanJS

OpenStudy (danjs):

Just have to fill in the numbers,, ,when you dot product two vectors together... jut add together the product of each dimensions coordinates.... v * w = 3*6 + 4*(-9) + 7*7 = ...

OpenStudy (usukidoll):

\[v \cdot w \] means take the dot product of vector v and vector w. nice explanation @DanJS :)